OpenStudy (anonymous):

2p^-5p-2=0 solve the equation by completing the square and leave the irrational roots in simplest radical form. please help solve.

OpenStudy (anonymous):

can you rewrite this?

OpenStudy (anonymous):

ya this isnt making much sense can you please rewrite this

OpenStudy (anonymous):

2p^2-5p-2=0 sorry.

OpenStudy (anonymous):

2p^2-5p=2 2(p^2-5/2p + 25/16)= 2 + 2(25/16) 2(p+5/4)^2=2+50/16 (p+5/4)^2=1+25/16 P=sqrt(1+25/16)-5/4 p= 0.350781059 and p= -2.85078106

OpenStudy (anonymous):

\[2p^2-5p-2=0\]\[p^2-\frac{5}{2}p=1\]\[p^2-\frac{5}{2}p+\frac{25}{16}=1+\frac{25}{16}\]\[\left(p-\frac{5}{4}\right)^2=\frac{41}{16}\]\[p-\frac{5}{4}=\pm\frac{\sqrt{41}}{4}\]\[p=\pm\frac{\sqrt{41}}{4}+\frac{5}{4}=\frac{5\pm\sqrt{41}}{4}\]